# Topology and Diffeology via Metric-like Functions

### v2

@@ 4. Uniformity defined by $d_{r}$
Since $K$ is compact, there exists $r^{\prime}>0$ such that $K\subset{r^{\prime}}D$. Setting $\varepsilon^{\prime}=\varepsilon$, we obtain the desired inclusion. Indeed, for any $H\in U_{r^{\prime},\varepsilon^{\prime}}[G]$, there exists $\varphi\colon U\to\phi(U)$ in $\mathcal{P}_{r^{\prime}}(G,H)$ such that $|p-\varphi(p)|<\varepsilon$ for all $p\in s(\varphi,r^{\prime})$. Since $\varphi\in\mathcal{P}_{r^{\prime}}(G,H)$, we have $K\cap G\subset{r^{\prime}}D\cap G\subset\varphi(U)$. Furthermore, for any $p\in K^{\varepsilon}\cap H\subset{r^{\prime}}D\cap H$, we obtain $|\varphi(p)-p|<\varepsilon^{\prime}=\varepsilon$, which implies that $\varphi(p)\in K$. Therefore, we conclude that $K^{\varepsilon}\cap H\subset\varphi^{-1}(K)$. Thus, $(U,\varphi(U),\varphi)$ is a triple $\varphi\colon H\dashrightarrow G$ that is $(\varepsilon,K)$-small, implying that $H\in\mathscr{U}_{K,\varepsilon}(G)$.
@@
Let $X$ be a set. A quasi-pseudo-metric on $X$ is a function $d\colon X\times X\to\mathbb{R}_{\geq 0}$ satisfying the following:
@@
The following argument is based on Proposition 2 in Chapter IX, Section 1.4 of [5] by Bourbaki.
@@
Using these $\{W_{n}\}_{n\in\mathbb{Z}_{\geq 0}}$, we define a function $g\colon X\times X\to\mathbb{R}_{\geq 0}$ by
$g(x,y)=\begin{cases}0&\text{if }(x,y)\in W_{n}\text{ for all }n\geq 1,\\ 2^{-k}&\text{if }(x,y)\in W_{n}\text{ for }1\leq n\leq k\text{ and }(x,y)\notin W_{k+1},\\ 1&\text{if }(x,y)\notin W_{1}.\end{cases}$
for $(x,y)\in X\times X$. Also, we define a function $d\colon X\times X\to\mathbb{R}_{\geq 0}$ by
@@
Suppose that the statement holds for every $p^{\prime}<p$. We now prove the case $p$. We define $a=\sum_{i=0}^{p-1}g(z_{i},z_{i+1})$. If $a\geq 1/2$, then the desired inequality holds trivially. If $a<1/2$, let $h$ be the largest integer $q\in\mathbb{Z}_{\geq 1}$ such that
$\sum_{i=1}^{q-1}g(z_{i},z_{i+1})\leq\dfrac{a}{2}.$
By the definition of $h$, we have
$\sum_{i=0}^{h-1}g(z_{i},z_{i+1})\leq\dfrac{a}{2}\quad\text{and}\quad\sum_{i=0}^{h}g(z_{i},z_{i+1})>\dfrac{a}{2}.$
It follows from the second inequality that $\sum_{i=h+1}^{p-1}g(z_{i},z_{i+1})\leq a/2$. By the inductive hypothesis, we have $g(x,z_{h})\leq a$ and $g(z_{h+1},y)\leq a$. It is also clear that $g(z_{h},z_{h+1})\leq a$. Let $k\in\mathbb{Z}_{\geq 1}$ be the smallest integer such that $2^{-k}\leq a$. Since $g(x,z_{h})$, $g(z_{h},z_{h+1})$, and $g(z_{h+1},y)$ are each of the form $2^{-n}$ and less than or equal to $a$, it follows that $(x,z_{h}),\ (z_{h},z_{h+1})$, and $(z_{h+1},y)$ all belong to $W_{k}$. Therefore, we obtain $(x,y)\in W_{k}\circ W_{k}\circ W_{k}\subset W_{k-1}$, which implies that $g(x,y)\leq 2^{1-k}\leq 2a$. This completes the proof of $\frac{1}{2}\,g(x,y)\leq d(x,y)$.
@@
If each $d_{r}$ is symmetric, then the family $\{d_{r}\}_{r>0}$ induces a genuine uniformity rather than just a quasi-uniformity. In this case, Proposition 2 in Chapter IX, Section 1.4 of [5] applies directly and ensures the existence of a pseudo-metric $d$ that induces the original uniformity. Moreover, if each $d_{r}$ is symmetric, then the topology induced by the family $\{d_{r}\}_{r>0}$ satisfies the $T_{1}$ separation axiom. In fact, it is even regular ([3], Lemma 4.3). As a consequence, the pseudo-metric $d$ constructed in the proof must actually be a metric.
@@
The family $\{d_{r}\}_{r>0}$ on the space of submanifolds $\Psi_{d}(\mathbb{R}^{N})$ in Example 2.7 is symmetric. Therefore, by Theorem 4.10 and Remark 4.11, there exists a metric on $\Psi_{d}(\mathbb{R}^{N})$ which induces both the uniformity and the topology on $\Psi_{d}(\mathbb{R}^{N})$. This provides a simpler way to recover the result of [3], Theorem 4.4.
@@ 5. Diffeology defined by $d_{r}$
In this section, we introduce a diffeology defined by the family $\{d_{r}\}_{r>0}$. We begin by defining diffeological spaces and see their basic properties. All definitions and properties presented here are based on [9] by Iglesias-Zemmour, the standard textbook of diffeology. For more details on diffeological spaces, see [9].
Definition 5.1 ([9], 1.5).
Let $X$ be a set. A parametrization of $X$ is a map $U\to X$ where $U$ is an open set of $\mathbb{R}^{n}$ for some $n\in\mathbb{Z}_{\geq 0}$. A diffeology on $X$ is a set $\mathscr{D}_{X}$ of parametrizations (whose elements are called plots) satisfying the following three axioms:
@@
Every constant parametrization $U\to X$ is a plot.
@@
Let $p\colon U\to X$ be a parametrization. If there exists an open covering $\{U_{\alpha}\}$ of $U$ such that ${p|}_{U_{\alpha}}\in\mathscr{D}_{X}$ for all $\alpha$, then $p$ itself is a plot.
@@
For every plot $p\colon U\to X$, every open set $V$ in $\mathbb{R}^{m}$ and $f\colon V\to U$ that is smooth as a map between Euclidean spaces, $p\circ f\colon V\to X$ is also a plot.
@@
Definition 5.2 ([9], 1.14).
Let $X$ and $Y$ be diffeological spaces. A map $f\colon X\to Y$ is said to be smooth if for every plot $p\colon U_{p}\to X$, the composition $f\circ p\colon U_{p}\to Y$ is a plot of $Y$. The set of smooth maps from $X$ to $Y$ is denoted by $C^{\infty}(X,Y)$. If $X$ is a Euclidean open set, then $C^{\infty}(X,Y)$ coincides with the set of all plots $X\to Y$ ([9], 1.16).
@@
It is known that the functor $D$ has the right adjoint $C$ (Shimakawa-Yoshida-Haraguchi [12], Proposition 2.1). For a manifold $M$ equipped with the standard diffeology (Example 5.3), the $D$-topology of $M$ coincides with the standard topology of the manifold $M$.
Definition 5.5 ([9], 1.66).
Let $X$ be a set, and let $\mathscr{F}$ be a family of parametrizations of $X$. Then, there exists the smallest diffeology $\mathscr{D}$ on $X$ that contains all parametrizations of $\mathscr{F}$ as plots. We call this diffeology $\mathscr{D}$ the diffeology generated by $\mathscr{F}$. A parametrization $p\colon U_{p}\to X$ is a plot of $\mathscr{D}$ if and only if it satisfies the following condition: there exists an open covering $\{U_{\alpha}\}$ of $U_{p}$ such that for each $\alpha$, either ${p|}_{U_{\alpha}}$ is a constant parametrization or ${p|}_{U_{\alpha}}=q\circ f$ for some $q\in\mathscr{F}$ and some smooth map $f\colon U_{\alpha}\to U_{q}$ ([9], 1.68).
If the family $\mathscr{F}$ satisfies $\bigcup_{q\in\mathscr{F}}q(U_{q})=X$, then $\mathscr{F}$ is called a covering generating family. In this case, a parametrization $p\colon U_{p}\to X$ is a plot of $\mathscr{D}$ if and only if it satisfies the following property: there exists an open covering $\{U_{\alpha}\}$ of $U_{p}$ such that for each $\alpha$, ${p|}_{U_{\alpha}}=q\circ f$ for some $q\in\mathscr{F}$ and some smooth map $f\colon U_{\alpha}\to U_{q}$.
@@
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$-axioms in Definition 4.4. A parametrization $p\colon U_{p}\to X$ is smooth at $t_{0}\in U_{p}$ if there exists $\delta>0$ satisfying the following conditions:
  • The map $B_{\delta}{(t_{0})}\to\mathbb{R};t\mapsto d_{r}(p(t_{0}),p(t))$ is continuous for each $r>0$.
  • For any $t_{1}\in B_{\delta}{(t_{0})}\setminus p^{-1}(p(t_{0}))$, the map $B_{\delta}{(t_{0})}\to\mathbb{R};t\mapsto d_{r}(p(t_{1}),p(t))$ is smooth at $t_{0}$ for each $r>0$.
A parametrization $p\colon U_{p}\to X$ is said to be a plot if it is smooth at every point $t_{0}\in U_{p}$. Then, the set of all plots of $X$ forms a diffeology on $X$.
@@
Since the covering and locality axioms follow directly from the definitions, we proceed to prove the smooth compatibility axiom. Let $p\colon U_{p}\to X$ be a plot of $X$, and let $f\colon V\to U_{p}$ be a smooth map from a Euclidean open set. Let $t_{0}\in V$ be arbitrary. Since $p$ is smooth at $f(t_{0})$, there exists $\delta>0$ such that the following conditions hold:
  • The map $B_{\delta}{(f(t_{0}))}\to\mathbb{R};t\mapsto d_{r}(p\circ f(t_{0}),p(t))$ is continuous for each $r>0$.
  • For any $t_{1}\in B_{\delta}{(f(t_{0}))}\setminus p^{-1}(p\circ f(t_{0}))$, the map $B_{\delta}{(f(t_{0}))}\to\mathbb{R};t\mapsto d_{r}(p\circ f(t_{1}),p(t))$ is smooth at $f(t_{0})$ for each $r>0$.
@@
$[t\mapsto d_{r}(p\circ f(s),p\circ f(t))]=[t\mapsto d_{r}(p\circ f(s),p(t))]\circ f$
for any $s\in B_{\delta^{\prime}}{(s)}$, we obtain:
  • The map $B_{\delta^{\prime}}{(t_{0})}\to\mathbb{R};t\mapsto d_{r}(p\circ f(t_{0}),p\circ f(t))$ is continuous.
  • For any $t_{1}\in B_{\delta^{\prime}}{(t_{0})}\setminus(p\circ f)^{-1}(p\circ f(t_{0}))$, the map $B_{\delta^{\prime}}{(t_{0})}\to\mathbb{R};t\mapsto d_{r}(p\circ f(t_{1}),p\circ f(t))$ is smooth at $t_{0}$ for each $r>0$.
@@
In [1], a diffeology is introduced on the tiling space. We aim to show that this diffeology coincides with the one defined using the family $\{d_{r}\}_{r>0}$ (Proposition 5.6). First, we review the diffeology on the tiling space defined in [1].
Definition 5.7 ([1], Definition 12).
@@
$F_{T_{1}}\colon\mathbb{R}^{d}\to\Omega_{T};x\mapsto T_{1}+x.$
The diffeology on $\Omega_{T}$ defined in [1] is the diffeology generated by the family $\{F_{T_{1}}\mid T_{1}\in\Omega_{T}\}$.
@@
First, we prove that $F_{T_{1}}$ is a plot of the diffeology induced by $\{d_{r}\}_{r>0}$ for each $T_{1}\in\Omega_{T}$. Let $\mathscr{P}$ be the set of prototiles of $T$, and let $T_{1}\in\Omega_{T}$ a tiling. Let $t_{0}\in\mathbb{R}^{d}$ be arbitrary, and define $\delta=\min_{\tau\in\mathscr{P}}(\mathop{\mathrm{diam}}{\tau})$. Then, for each $t\in B_{\delta}{(t_{0})}$, we have
@@
This holds because $d_{r}(T_{1}+t_{0},T_{1}+t)$ cannot be smaller than $\delta=\min_{\tau\in\mathscr{P}}(\mathop{\mathrm{diam}}{\tau})$. Thus, $F_{T_{1}}$ satisfies the conditions specified in Proposition 5.6 for each $T_{1}\in\Omega_{T}$.
Next, we prove the other inclusion. Let $p\colon U_{p}\to\Omega_{T}$ be a plot in the sense of Proposition 5.6, and let $t_{0}\in U_{p}$. Then, there exists $\delta>0$ such that the map
$B_{\delta}{(t_{0})}\to\mathbb{R};t\mapsto d_{r}(p(t_{0}),p(t))$
is continuous for each $r>0$. Define $\varepsilon>0$ such that $0<\varepsilon<1/2\min_{\tau\in\mathscr{P}}(\mathop{\mathrm{diam}}{\tau})$. Let $r>0$ be arbitrary. By the continuity of the map $t\mapsto d_{r}(p(t_{0}),p(t))$, we can choose $\delta>0$ such that $d_{r}(p(t_{0}),p(t))<\varepsilon$ for all $t\in B_{\delta}{(t_{0})}$. In fact, this $\delta$ satisfies $d_{r^{\prime}}(p(t_{0}),p(t))<\varepsilon$ for all $r^{\prime}>0$ and all $t\in B_{\delta}{(t_{0})}$. To prove the above statement, it suffices to show that the set
$A_{r^{\prime}}=\{t\in B_{\delta}{(t_{0})}\mid d_{r^{\prime}}(p(t_{0}),p(t))<\varepsilon\}\,(\neq\emptyset)$
is connected in $B_{\delta}{(t_{0})}$ for all $r^{\prime}>0$. We only need to consider the case where $r^{\prime}>r$, as $r^{\prime}\leq r$ implies $A_{r^{\prime}}\supset A_{r}=B_{\delta}{(t_{0})}$. Since the map $d_{r^{\prime}}(p(t_{0}),\cdot)\colon X\to\mathbb{R}$ is upper semi-continuous (Corollary 2.5), $A_{r^{\prime}}\subset B_{\delta}{(t_{0})}$ is open. Therefore, it suffices to show that $A_{r^{\prime}}$ is closed in $B_{\delta}{(t_{0})}$. To prove this, suppose there exists $t\in\overline{A_{r^{\prime}}}\setminus A_{r^{\prime}}$. Since $t\notin A_{r^{\prime}}$, we have $d_{r^{\prime}}(p(t_{0}),p(t))\geq\varepsilon$. On the other hand, since $t\in B_{\delta}{(t_{0})}$, there exists $v\in\mathbb{R}^{d}$ such that $|v|<\varepsilon$ and

### v3

@@ 4. Uniformity defined by $d_{r}$
Since $K$ is compact, there exists $r^{\prime}>0$ such that $K\subset{r^{\prime}}D$. Setting $\varepsilon^{\prime}=\varepsilon$, we obtain the desired inclusion. Indeed, for any $H\in U_{r^{\prime},\varepsilon^{\prime}}[G]$, there exists $\varphi\colon U\rightarrow\phi(U)$ in $\mathcal{P}_{r^{\prime}}(G,H)$ such that $|p-\varphi(p)|<\varepsilon$ for all $p\in s(\varphi,r^{\prime})$. Since $\varphi\in\mathcal{P}_{r^{\prime}}(G,H)$, we have $K\cap G\subset{r^{\prime}}D\cap G\subset\varphi(U)$. Furthermore, for any $p\in K^{\varepsilon}\cap H\subset{r^{\prime}}D\cap H$, we obtain $|\varphi(p)-p|<\varepsilon^{\prime}=\varepsilon$, which implies that $\varphi(p)\in K$. Therefore, we conclude that $K^{\varepsilon}\cap H\subset\varphi^{-1}(K)$. Thus, $(U,\varphi(U),\varphi)$ is a triple $\varphi\colon H\dashrightarrow G$ that is $(\varepsilon,K)$-small, implying that $H\in\mathscr{U}_{K,\varepsilon}(G)$.
@@
Let $X$ be a set. A quasi-pseudo-metric on $X$ is a function $d\colon X\times X\rightarrow\mathbb{R}_{\geq 0}$ satisfying the following:
@@
The following argument is based on Proposition 2 in Chapter IX, Section 1.4 of [MR1726872] by Bourbaki.
@@
Using these $\{W_{n}\}_{n\in\mathbb{Z}_{\geq 0}}$, we define a function $g\colon X\times X\rightarrow\mathbb{R}_{\geq 0}$ by
[…]
for $(x,y)\in X\times X$. Also, we define a function $d\colon X\times X\rightarrow\mathbb{R}_{\geq 0}$ by
@@
The case $p=1$ is immediate. Suppose that the statement holds for every $p^{\prime}<p$. We now prove the case $p$. We define $a=\sum_{i=0}^{p-1}g(z_{i},z_{i+1})$. If $a\geq 1/2$, then the desired inequality holds trivially. If $a=0$, the claim is clear. Thus assume $0<a<1/2$.
If there exists $q\in\{1,\dots,p-1\}$ such that
$\sum_{i=0}^{q-1}g(z_{i},z_{i+1})\leq\dfrac{a}{2},$
let $h$ be the largest such integer; otherwise, set $h=0$. Then
$\sum_{i=0}^{h-1}g(z_{i},z_{i+1})\leq\dfrac{a}{2}\quad\text{and}\quad\sum_{i=0}^{h}g(z_{i},z_{i+1})>\dfrac{a}{2},$
where the first sum is understood to be 0 when $h=0$. It follows that $\sum_{i=h+1}^{p-1}g(z_{i},z_{i+1})<a/2$.
Applying the inductive hypothesis to each nontrivial one of the two subsequences $z_{0},\dots,z_{h}$ and $z_{h+1},\dots,z_{p}$ (and using $g(x,x)=g(y,y)=0$ in the trivial case), we obtain $g(x,z_{h})\leq a$ and $g(z_{h+1},y)\leq a$. It is also clear that $g(z_{h},z_{h+1})\leq a$.
Let $k\in\mathbb{Z}_{\geq 1}$ be the smallest integer such that $2^{-k}\leq a$. By the definition of $g$, it follows that $(x,z_{h}),\ (z_{h},z_{h+1})$, and $(z_{h+1},y)$ all belong to $W_{k}$. Therefore, we obtain
$(x,y)\in W_{k}\circ W_{k}\circ W_{k}\subset W_{k-1},$
which implies that $g(x,y)\leq 2^{1-k}\leq 2a$.
@@
If each $d_{r}$ is symmetric, then the family $\{d_{r}\}_{r>0}$ induces a genuine uniformity rather than just a quasi-uniformity. In this case, Proposition 2 in Chapter IX, Section 1.4 of [MR1726872] applies directly and ensures the existence of a pseudo-metric $d$ that induces the original uniformity. Moreover, if each $d_{r}$ is symmetric, then the topology induced by the family $\{d_{r}\}_{r>0}$ satisfies the $T_{1}$ separation axiom. In fact, it is even regular ([MR3243393], Lemma 4.3). As a consequence, the pseudo-metric $d$ constructed in the proof must actually be a metric.
@@
The family $\{d_{r}\}_{r>0}$ on the space of submanifolds $\Psi_{d}(\mathbb{R}^{N})$ in Example 2.7 is symmetric. Therefore, by Theorem 4.10 and Remark 4.11, there exists a metric on $\Psi_{d}(\mathbb{R}^{N})$ which induces both the uniformity and the topology on $\Psi_{d}(\mathbb{R}^{N})$. This provides a simpler way to recover the result of [MR3243393], Theorem 4.4.
@@ 5. Diffeology defined by $d_{r}$
In this section, we introduce a diffeology defined by the family $\{d_{r}\}_{r>0}$. We begin by defining diffeological spaces and see their basic properties. All definitions and properties presented here are based on [MR3025051] by Iglesias-Zemmour, the standard textbook of diffeology. For more details on diffeological spaces, see [MR3025051].
Definition 5.1 ([MR3025051], 1.5).
Let $X$ be a set. A parametrization of $X$ is a map $U\rightarrow X$ where $U$ is an open set of $\mathbb{R}^{n}$ for some $n\in\mathbb{Z}_{\geq 0}$. A diffeology on $X$ is a set $\mathscr{D}_{X}$ of parametrizations (whose elements are called plots) satisfying the following three axioms:
@@
Every constant parametrization $U\rightarrow X$ is a plot.
@@
Let $p\colon U\rightarrow X$ be a parametrization. If there exists an open covering $\{U_{\alpha}\}$ of $U$ such that ${p|}_{U_{\alpha}}\in\mathscr{D}_{X}$ for all $\alpha$, then $p$ itself is a plot.
@@
For every plot $p\colon U\rightarrow X$, every open set $V$ in $\mathbb{R}^{m}$ and $f\colon V\rightarrow U$ that is smooth as a map between Euclidean spaces, $p\circ f\colon V\rightarrow X$ is also a plot.
@@
Definition 5.2 ([MR3025051], 1.14).
Let $X$ and $Y$ be diffeological spaces. A map $f\colon X\rightarrow Y$ is said to be smooth if for every plot $p\colon U_{p}\rightarrow X$, the composition $f\circ p\colon U_{p}\rightarrow Y$ is a plot of $Y$. The set of smooth maps from $X$ to $Y$ is denoted by $C^{\infty}(X,Y)$. If $X$ is a Euclidean open set, then $C^{\infty}(X,Y)$ coincides with the set of all plots $X\rightarrow Y$ ([MR3025051], 1.16).
@@
It is known that the functor $D$ has the right adjoint $C$ (Shimakawa-Yoshida-Haraguchi [MR3884529], Proposition 2.1). For a manifold $M$ equipped with the standard diffeology (Example 5.3), the $D$-topology of $M$ coincides with the standard topology of the manifold $M$.
Definition 5.5 ([MR3025051], 1.66).
Let $X$ be a set, and let $\mathscr{F}$ be a family of parametrizations of $X$. Then, there exists the smallest diffeology $\mathscr{D}$ on $X$ that contains all parametrizations of $\mathscr{F}$ as plots. We call this diffeology $\mathscr{D}$ the diffeology generated by $\mathscr{F}$. A parametrization $p\colon U_{p}\rightarrow X$ is a plot of $\mathscr{D}$ if and only if it satisfies the following condition: there exists an open covering $\{U_{\alpha}\}$ of $U_{p}$ such that for each $\alpha$, either ${p|}_{U_{\alpha}}$ is a constant parametrization or ${p|}_{U_{\alpha}}=q\circ f$ for some $q\in\mathscr{F}$ and some smooth map $f\colon U_{\alpha}\rightarrow U_{q}$ ([MR3025051], 1.68).
If the family $\mathscr{F}$ satisfies $\bigcup_{q\in\mathscr{F}}q(U_{q})=X$, then $\mathscr{F}$ is called a covering generating family. In this case, a parametrization $p\colon U_{p}\rightarrow X$ is a plot of $\mathscr{D}$ if and only if it satisfies the following property: there exists an open covering $\{U_{\alpha}\}$ of $U_{p}$ such that for each $\alpha$, ${p|}_{U_{\alpha}}=q\circ f$ for some $q\in\mathscr{F}$ and some smooth map $f\colon U_{\alpha}\rightarrow U_{q}$.
@@
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$-axioms in Definition 4.4. A parametrization $p\colon U_{p}\rightarrow X$ is smooth at $t_{0}\in U_{p}$ if there exists $\delta>0$ satisfying the following conditions:
  • The map $B_{\delta}{(t_{0})}\rightarrow\mathbb{R};t\mapstochar\rightarrow d_{r}(p(t_{0}),p(t))$ is continuous for each $r>0$.
  • For any $t_{1}\in B_{\delta}{(t_{0})}\setminus p^{-1}(p(t_{0}))$, the map $B_{\delta}{(t_{0})}\rightarrow\mathbb{R};t\mapstochar\rightarrow d_{r}(p(t_{1}),p(t))$ is smooth at $t_{0}$ for each $r>0$.
A parametrization $p\colon U_{p}\rightarrow X$ is said to be a plot if it is smooth at every point $t_{0}\in U_{p}$. Then, the set of all plots of $X$ forms a diffeology on $X$.
@@
Since the covering and locality axioms follow directly from the definitions, we proceed to prove the smooth compatibility axiom. Let $p\colon U_{p}\rightarrow X$ be a plot of $X$, and let $f\colon V\rightarrow U_{p}$ be a smooth map from a Euclidean open set. Let $t_{0}\in V$ be arbitrary. Since $p$ is smooth at $f(t_{0})$, there exists $\delta>0$ such that the following conditions hold:
  • The map $B_{\delta}{(f(t_{0}))}\rightarrow\mathbb{R};t\mapstochar\rightarrow d_{r}(p\circ f(t_{0}),p(t))$ is continuous for each $r>0$.
  • For any $t_{1}\in B_{\delta}{(f(t_{0}))}\setminus p^{-1}(p\circ f(t_{0}))$, the map $B_{\delta}{(f(t_{0}))}\rightarrow\mathbb{R};t\mapstochar\rightarrow d_{r}(p\circ f(t_{1}),p(t))$ is smooth at $f(t_{0})$ for each $r>0$.
@@
$t\mapstochar\rightarrow d_{r}(p\circ f(s),p\circ f(t))]=[t\mapstochar\rightarrow d_{r}(p\circ f(s),p(t))]\circ f$
for any $s\in B_{\delta^{\prime}}{(t_{0})}$, we obtain:
  • The map $B_{\delta^{\prime}}{(t_{0})}\rightarrow\mathbb{R};t\mapstochar\rightarrow d_{r}(p\circ f(t_{0}),p\circ f(t))$ is continuous.
  • For any $t_{1}\in B_{\delta^{\prime}}{(t_{0})}\setminus(p\circ f)^{-1}(p\circ f(t_{0}))$, the map $B_{\delta^{\prime}}{(t_{0})}\rightarrow\mathbb{R};t\mapstochar\rightarrow d_{r}(p\circ f(t_{1}),p\circ f(t))$ is smooth at $t_{0}$ for each $r>0$.
@@
In [MR4882913], a diffeology is introduced on the tiling space. We aim to show that this diffeology coincides with the one defined using the family $\{d_{r}\}_{r>0}$ (Proposition 5.6). First, we review the diffeology on the tiling space defined in [MR4882913].
Definition 5.7 ([MR4882913], Definition 12).
@@
$F_{T_{1}}\colon\mathbb{R}^{d}\rightarrow\Omega_{T};x\mapstochar\rightarrow T_{1}+x.$
The diffeology on $\Omega_{T}$ defined in [MR4882913] is the diffeology generated by the family $\{F_{T_{1}}\mid T_{1}\in\Omega_{T}\}$.
@@
Let $\mathscr{P}$ be the set of prototiles of $T$. For each $\tau\in\mathscr{P}$, choose $\lambda_{\tau}>0$ such that
$\operatorname{Int}{(}\tau)\cap\operatorname{Int}{(}\tau+v)\neq\emptyset$
for every $v\in\mathbb{R}^{d}$ with $0<|v|<2\lambda_{\tau}$, and set
$\lambda_{T}:=\min_{\tau\in\mathscr{P}}\lambda_{\tau}>0.$
Then, for every $S\in\Omega_{T}$, every nonempty subset $K\subset\mathbb{R}^{d}$, and any $u,v\in\mathbb{R}^{d}$ with $|u|,|v|<\lambda_{T}$, if $(S+u)\cap K=(S+v)\cap K$, then $u=v$.
First, we prove that $F_{T_{1}}$ is a plot of the diffeology induced by $\{d_{r}\}_{r>0}$ for each $T_{1}\in\Omega_{T}$. Let $T_{1}\in\Omega_{T}$ be a tiling, and let $t_{0}\in\mathbb{R}^{d}$ be arbitrary. Set $\delta=\lambda_{T}/2$. Then, for each $t\in B_{\delta}{(t_{0})}$, we have
@@
Indeed, since $(T_{1}+t_{0})+(t-t_{0})=T_{1}+t$, we have $d_{r}(T_{1}+t_{0},T_{1}+t)\leq|t_{0}-t|$. If strict inequality held, then there would exist $v\in\mathbb{R}^{d}$ with $|v|<|t_{0}-t|<\lambda_{T}$ such that $((T_{1}+t_{0})+v)\cap rD=(T_{1}+t)\cap rD$. Since also $(T_{1}+t_{0})+(t-t_{0})=T_{1}+t$ and $|t-t_{0}|<\lambda_{T}$, the defining property of $\lambda_{T}$ implies $v=t-t_{0}$, a contradiction. Moreover, for any $t_{1}\in B_{\delta}{(t_{0})}$, the same argument shows that $d_{r}(F_{T_{1}}(t_{1}),F_{T_{1}}(t))=|t_{1}-t|$ for each $t\in B_{\delta}{(t_{0})}$, since $t,t_{1}\in B_{\delta}{(t_{0})}$ implies $|t_{1}-t|<2\delta=\lambda_{T}$. In particular, the map $t\mapstochar\rightarrow d_{r}(F_{T_{1}}(t_{1}),F_{T_{1}}(t))=|t_{1}-t|$ is smooth at $t_{0}$ for each $t_{1}\neq t_{0}$. Thus, $F_{T_{1}}$ satisfies the conditions specified in Proposition 5.6 for each $T_{1}\in\Omega_{T}$.
Next, we prove the other inclusion. Let $p\colon U_{p}\rightarrow\Omega_{T}$ be a plot in the sense of Proposition 5.6, and let $t_{0}\in U_{p}$. Then, there exists $\delta>0$ such that the map
$B_{\delta}{(t_{0})}\rightarrow\mathbb{R};\ t\mapstochar\rightarrow d_{r}(p(t_{0}),p(t))$
is continuous for each $r>0$. Choose $\varepsilon>0$ such that $0<\varepsilon<\lambda_{T}/2$. Let $r>0$ be arbitrary. By the continuity of the map $t\mapstochar\rightarrow d_{r}(p(t_{0}),p(t))$, we can choose $0<\delta^{\prime}\leq\delta$ such that $d_{r}(p(t_{0}),p(t))<\varepsilon$ for all $t\in B_{\delta^{\prime}}{(t_{0})}$. We claim that this $\delta^{\prime}$ satisfies $d_{r^{\prime}}(p(t_{0}),p(t))<\varepsilon$ for all $r^{\prime}>0$ and all $t\in B_{\delta^{\prime}}{(t_{0})}$. To prove this, define
$A_{r^{\prime}}=\{t\in B_{\delta^{\prime}}{(t_{0})}\mid d_{r^{\prime}}(p(t_{0}),p(t))<\varepsilon\}.$
We only need to consider the case where $r^{\prime}>r$, as $r^{\prime}\leq r$ implies $A_{r^{\prime}}\supset A_{r}=B_{\delta^{\prime}}{(t_{0})}$. Since the map $t\mapstochar\rightarrow d_{r^{\prime}}(p(t_{0}),p(t))$ is continuous on $B_{\delta}{(t_{0})}$, $A_{r^{\prime}}\subset B_{\delta^{\prime}}{(t_{0})}$ is open. Since $B_{\delta^{\prime}}{(t_{0})}$ is connected and $t_{0}\in A_{r^{\prime}}$, it suffices to show that $A_{r^{\prime}}$ is closed in $B_{\delta^{\prime}}{(t_{0})}$. To prove this, suppose there exists $t\in\overline{A_{r^{\prime}}}\setminus A_{r^{\prime}}$. Since $t\in B_{\delta^{\prime}}{(t_{0})}$, there exists $v\in\mathbb{R}^{d}$ such that $|v|<\varepsilon$ and